OpenStudy (anonymous):

We toss two dice 1,000 times. How many times do we expect to have the sum of the two dice equal to 4?

OpenStudy (zarkon):

do you know the probability if you roll two die that the sum is 4?

OpenStudy (anonymous):

no

OpenStudy (p0sitr0n):

P(4)=(1+3)+(3+1)+(2+2) / 36= 3/36=1/12 This is the probability density. Now, N*P= 1/12 * 1000= around 83.33 times